Accessibility Assistance

Skip to Content




Information Technology
George A. Smathers Libraries
University of Florida
Suite 1250
2124 NE Waldo Rd
Gainesville, FL 32609

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
July 27, 2010

Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Winston Harris, Michael Jay, Tom Minton, Chris Nicolich, Patricia Ruwell, Josh Spurgin, Evan Wack
Announcements
  1. Winston has purchased some real forks and plates for the Eastside IT kitchen. If used, please wash.
  2. TSS meeting was cancelled last week. Please review the TSS brainstorming topics and email Will with the list of topics in which you are interested in participating. Include the degree of participation.
  3. Library Council - changes in health care benefits discussed. For further information, check Info Gator.
    Library budget - flat budget. OPS is the same as last year.
    Merit raises - retroactive as of July 1st for those who ranked achieves as well as exceeds. Percentages not yet determined. Rate change should appear in the August 27 paycheck.
    ARL stats displayed: pat on the back to Tom Bielicke, Winston and Michael.
    Newell Hall project - should be vacant in the first quarter of 2011. Date of completion unknown.
    Library wishlist - for first-time donors to contribute.
    LibGuides - are now embedded in Sakai.
    Monday, August 2, 1:30-2:30PM - mandatory Town Hall meeting on Open Access in Room 1A.
    Parking pass for campus is kept in Will's office. It is available for all to use.
  4. Eastside front door to building now set for 6:45AM.
  5. If you need shelving for your office, please submit a Facilities request for the installation of the leftover West shelving, currently housed at ALF.
Review Minutes of 07/06/10
  1. SW #2 - ASERL not ASURL.
Ongoing IT Updates

Hardware
  1. Public printing - is now back up everywhere. There were a couple of rogue machines with a password protection error. Josh will be training with Matt. Outage was due to changes on the server to enable laptop printing. Printing was down over the weekend. There is a plan in place to roll out new drivers. Printing now works from checkout laptops. Duplex is working. In three weeks, Canon will certify new software to fix all problems. Lag time has been fixed - Arrowmark designed a better gateway. Yes, patrons will be able to print from their own personal laptops. We will need to put a link for it on the printing solution page. Barbara Hood will need to publicize. Shoot an email to Will so he can get Admin to decide.
  2. 1A AC is fixed. It is now 90% quieter. Now we need to arrange for the AV team to come out and reconfigure the ceiling mikes.
  3. Network configuration problems - logging on key server. Matt Grover is working on it.
  4. Public imaging - finishing microforms. Music and Journalism are done. AFA will be done this week, and then Education and MSL to follow. August 9 - West.
  5. August 9-13 will be the week all of IT pitches in to take out 138 old machines in West and install the new ones. We will need to rerun all cables and monitors and set the machines to PXE boot. All this, assuming the weird networking problem is fixed.
  6. Preservation move - six chairs from our old area will be brought out to the Eastside workshop.
  7. Privilege Authority - given to the VIVO staff.
  8. Staff migration on hold for public.
  9. Time Limited workstations - ready to demo on Monday. How to stop people from going to other computers? Amy is working on that issue. Gatorlinks need to be restricted to those five machines only. Logan and Amy need to meet with Public Services folks to work out the details.
  10. Help Desk - outage update - need a procedure to get the word out via phone announcement.
  11. Logan notes we are running out of space on the Z & U drives on the SAN. We need to work on a solution, even if it is to add more space.
Software
  1. Open Access Publication Fund form is done.
  2. Document disposition database - Winston will be accompanying Jan, Judy and Chelsea to ASERL in Atlanta. Will determine if new requirements are needed.
  3. Database migration - Chris has made good progress. 2K3SQL and SQL08 can be decommissioned.
  4. Logan created a new server for ILL: LIB-ILL. Need to download third party tool "Always Up" for ARIEL.
  5. FCLA Aleph report - one is in the works.
  6. Tom Minton identified a site for a library hours page at the University of Michigan. Need to come up with components for our Library main page.
  7. SAN applications folder - 80-90% of all apps have been moved. There is a problem with Nelda's machine - she can't connect.
  8. Chris has been working with Evan's group on tools to work with Delphi apps. Label printing, etc - all tools for staff Win7 migration.
  9. Macros - Acquisitions is ongoing; Michael is working with Paul Lightcap. Panama Canal items macro - done. Minor Cataloging macro - done. Working on a difficult macro left behind by Jason Fleming. Analyzing a request from Paul.
  10. The OLE candidate accepted: Maureen Kelly will start on August 27.
Network/Server Administration
  1. Sharepoint - Logan is working on this and will need some direction for it. Cataloging and Acquisitions have already expressed a desire for it. Training? Will is looking into New Horizons. We can crank out a basic look where people can drop in their own components. Right now, Logan is the only administrator. Winston suggests that the OPS documents on the Fiscal Services website would be a good candidate for a Sharepoint site. Cynthia would like to see the public folders migrated to a Sharepoint site.
  2. Virtualization - VIVO is pretty much done. There were a few hiccups, but we got through it. They want to maintain test boxes. App server to be done the week after next (when Chris returns from vacation). Staff printing is ready to go - will wait for Amy to return from vacation. GIS - in the process of moving. Should be done by the end of the week. Physical move of MSL server to CSE in about 2-3 weeks. Need to move UFDC data on Web 2 then we can take MSL down.
  3. Altiris server move to the physical server - by the end of next week. Need to put in a request for new drives for that server.
ILS
  1. SQL queries - Cynthia has been working with Donna Alsbury at FCLA. Some of the feedback that Cynthia has received is that the reports are not accurate.
  2. Cynthia is starting to get her SQL query mojo back.
  3. Some of the tickets are not clear - Steve, Raimonda and Betsy are being quite patient.
  4. Matthew notes that some of the people who were tapped to take over aspects of Jason's job when he left were not officially told of their new duties. Matthew will put together an email to this effect and send it to Will.
DLC General Issues
  1. Migration is just about done. All working correctly.
  2. OCR - working on issues. ABBY - problems for VM. Need drivers and dongle installed on a machine in DLC. Will get the dongle back from CNS and give it to DLC. That should solve the problem.
Web Projects
  1. Matthew is working on a video for Rachel Schipper.
  2. Matthew has been redesigning the Judaica site.
  3. Tom has been reorganizing the Fiscal Services site.
  4. Proxy server - need to get the groups straightened out. May need to create a specific Library group or a different group just for the proxy. Cliff Richmond to look into it. Proxy issues: failover, authentication, and setting it up to recognize off-campus, VPN and log in.
  5. Word Press - DOA. Judy did not care for the design. Back to the drawing board. CMS still in the mix.
  6. Web server to migrate to virtual server as is.